# Information-Theoretic Limits of Bistatic Integrated Sensing and Communication

### v3

@@ Appendix C Proof of Theorem 5
$\quad\ nR-n\epsilon_{n}\leq I\left(W;Y^{n},S^{n}\right)=I\left(W;Y^{n}|S^{n}\right)=\sum_{i=1}^{n}I\left(W;Y_{i}|Y^{i-1},S^{n}\right)=\sum_{i=1}^{n}H\left(Y_{i}|Y^{i-1},S^{n}\right)$
$-H\left(Y_{i}|X_{i},W,Y^{i-1},S^{n}\right)\stackrel{{\scriptstyle(a)}}{{\leq}}\sum_{i=1}^{n}H\left(Y_{i}|S_{i}\right)-H\left(Y_{i}|X_{i},S_{i}\right)=\sum_{i=1}^{n}I(X_{i};Y_{i}|S_{i}),$ (29)
@@
$\quad\ nR-n\epsilon_{n}\leq I\left(W;Y^{n},S^{n}\right)\leq\sum_{i=1}^{n}I\left(X,Z_{i+1}^{n};Y_{i},S_{i}|Y^{i-1},S^{i-1}\right)+I\left(Z_{i+1}^{n};Z_{i}\right)$
$\leq\sum_{i=1}^{n}I\left(Z_{i+1}^{n};Y_{i},S_{i}|Y^{i-1},S^{i-1}\right)+I\left(X;Y_{i},S_{i}|Y^{i-1},S^{i-1},Z_{i+1}^{n}\right)+I\left(Z_{i+1}^{n};Z_{i}\right)$
=$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{i=1}^{n}I\left(Y^{i-1},S^{i-1};Z_{i}|Z_{i+1}^{n}\right)+I\left(X;Y_{i},S_{i}|Y^{i-1},S^{i-1},Z_{i+1}^{n}\right)+I\left(Z_{i+1}^{n};Z_{i}\right)$
@@
$=\sum_{i=1}^{n}I\left(X_{i};Y_{i},Z_{i},S_{i}|U_{i},V_{i}\right)+I\left(U_{i},V_{i};Z_{i}\right)=\sum_{i=1}^{n}I\left(X_{i};Y_{i},Z_{i}|S_{i},U_{i},V_{i}\right)+I\left(U_{i},V_{i};Z_{i}\right)$ (31)
The rest of the proof of the rate expression follows by introducing a time-sharing random variable $Q\sim$ Unif $[1:n]$ independent of $\left(W,X^{n},Y^{n},Z^{n},S^{n}\right)$ and defining $U=\left(Q,U_{Q}\right),V=\left(Q,V_{Q}\right),X=X_{Q},Y=Y_{Q}$, $Z=Z_{Q}$ and $S=S_{Q}$.
@@ Appendix D Proof of Theorem 7
The proof of the distortion constraint is similar to that in Theorem 5, which completes the proof of the converse.
Appendix E Bounds on the capacity-distortion function of Example 2
E-A Three lower bounds to Example 2
@@
$I(X;Y|S)$ $=p_{S}(0)H(Y|S=0)+p_{S}(1)H(Y|S=1)=p_{S}(1)H(X)=qH_{\mathrm{2}}(p).$
Notice that the estimator SenRx accurately estimate the state $\hat{S}=S=1$ when $z=a$ or $z=a+1$. Since the distortion occurs at $z=0$ and $z=1$, we then calculate the distortion in two cases. For $z=0$, we need to compare $p(s=0|z=0)$ and $p(s=1|z=0)$ and select the corresponding $S$ with a higher probability as the estimate of $S$. And we can calculate that
@@
Applying Theorem 3 to Example 2, since $\left|\mathcal{U}\right|\leq|\mathcal{X}|$, without loss of generality, we take $\mathcal{U}=\{0,1\}$, $u\in\mathcal{U}$, and let the joint distribution of $U$ and $X$ be $p_{jk}=p(u=j,x=k),j=0,1,k=0,1$ and define $p_{uj}=p(u=j),j=0,1$. Next, we have
$I\left(X;Y|U,S\right)=qH(Y|U,S=1)=q\Big{(}p_{u0}H_{2}\big{(}\frac{p_{00}}{p_{u0}}\big{)}+p_{u1}H_{2}\big{(}\frac{p_{10}}{p_{u1}}\big{)}\Big{)}.$
According to $\hat{s}^{*}(u,z)$ in Theorem 3, for $z=0$, we can calculate that
@@
$C(D)$ $\geq\max_{p_{UX}}\{\min\{qH_{2}(p),I\left(U;Z\right)+I\left(X;Y|U,S\right)\}:\sum_{i=0}^{1}\sum_{j=0}^{1}p(z=i,u=j)d_{ij}\leq D\}$
@@
Applying Corollary 2 to Example 2, we can get the optimal estimator for $z=0$ and $z=1$,
@@
where we used the independence of $S$ and $X$. Thus, the expected distortion of the optimal estimator is
$D=\sum_{x}P_{X}(x)c(x)=P_{X}(0)c(0)=p\min\{q,1-q\},$
Therefore, we obtain
$C(D)$ $=\max_{p\in[0,1]}\{\min\{qH_{2}(p),I(X;Z)\}:p\min\{q,1-q\}\leq D\}.$
E-B Two upper bounds to Example 2

### v4

@@ Appendix C Proof of Theorem 5
$\quad\ nR-n\epsilon_{n}\leq I\left(W;Y^{n},S^{n}\right)=I\left(W;Y^{n}|S^{n}\right)$
$=\sum_{i=1}^{n}I\left(W;Y_{i}|Y^{i-1},S^{n}\right)=\sum_{i=1}^{n}H\left(Y_{i}|Y^{i-1},S^{n}\right)-H\left(Y_{i}|X_{i},W,Y^{i-1},S^{n}\right)$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\sum_{i=1}^{n}H\left(Y_{i}|S_{i}\right)-H\left(Y_{i}|X_{i},S_{i}\right)=\sum_{i=1}^{n}I(X_{i};Y_{i}|S_{i}),$ (27)
@@
$\quad\ nR-n\epsilon_{n}\leq I\left(W;Y^{n},S^{n}\right)$
$\leq\sum_{i=1}^{n}I\left(W,Z_{i+1}^{n};Y_{i},S_{i}|Y^{i-1},S^{i-1}\right)+I\left(Z_{i+1}^{n};Z_{i}\right)$
$=\sum_{i=1}^{n}I\left(Z_{i+1}^{n};Y_{i},S_{i}|Y^{i-1},S^{i-1}\right)+I\left(W;Y_{i},S_{i}|Y^{i-1},S^{i-1},Z_{i+1}^{n}\right)+I\left(Z_{i+1}^{n};Z_{i}\right)$
=$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{i=1}^{n}I\left(Y^{i-1},S^{i-1};Z_{i}|Z_{i+1}^{n}\right)+I\left(W;Y_{i},S_{i}|Y^{i-1},S^{i-1},Z_{i+1}^{n}\right)+I\left(Z_{i+1}^{n};Z_{i}\right)$
@@
$=\sum_{i=1}^{n}I\left(X_{i};Y_{i},Z_{i},S_{i}|U_{i},V_{i}\right)+I\left(U_{i},V_{i};Z_{i}\right)$
$=\sum_{i=1}^{n}I\left(X_{i};Y_{i},Z_{i}|S_{i},U_{i},V_{i}\right)+I\left(U_{i},V_{i};Z_{i}\right)$ (29)
The rest of the proof of the rate expression follows by exploiting the time-sharing random variable $Q\sim$ Unif $[1:n]$ introduced in the proof of Theorem 5. The cardinality bounds on $U$ and $V$ can be proved using the extension of the convex cover method to multiple random variables in [45, Appendix C].
@@ Appendix D Proof of Theorem 8
The proof of the distortion constraint is identical to that in Theorem 5 and is omitted here.
Appendix E Proof of Corollary 5
and
$\sum_{i=1}^{n}\mathbb{E}[d(\hat{S}_{i},S_{i})]\leq nD.$
Using Fano’s inequality, we have that
$\leq$ $\frac{1}{n}\sum_{i=1}^{n}H(X_{i}|\hat{X}^{n})\leq\frac{1}{n}\sum_{i=1}^{n}H(X_{i}|\hat{X}_{i})\leq\frac{1}{n}\sum_{i=1}^{n}{\rm{H}}_{2}(p_{i})+p_{i}\log(|\mathcal{X}|-1)$
$\leq$ ${\rm{H}}_{2}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}p_{i}\bigg{)}={\rm{H}}_{2}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[d(\hat{S}_{i},S_{i})]\bigg{)}.$ (32)
Hence we obtain
$\frac{1}{n}H(X^{n})-\frac{1}{n}I(X^{n};Z^{n})\leq{\rm{H}}_{2}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[d(\hat{S}_{i},S_{i})]\bigg{)}.$ (33)
Combining the fact that $H(X^{n})=nR$ with $\frac{1}{n}I(X^{n};Z^{n})\leq 1-{\rm{H}}_{2}(q)$, we get ${\rm{H}}_{2}^{-1}\big{(}R-1+{\rm{H}}_{2}(q)\big{)}\leq D$, which completes the proof.
Appendix F Bounds on the capacity-distortion function of Example 2
F-A Three lower bounds to Example 2
@@
$I(X;Y|S)$ $=p_{S}(0)H(Y|S=0)+p_{S}(1)H(Y|S=1)=p_{S}(1)H(X)=q{\rm{H}}_{2}(p).$
Notice that the estimator at the SenRx accurately estimate the state $\hat{S}=S=1$ when $z=a$ or $z=a+1$. Since the distortion occurs at $z=0$ and $z=1$, we then calculate the distortion in two cases. For $z=0$, we need to compare $p(s=0|z=0)$ and $p(s=1|z=0)$ and select the corresponding $S$ with a higher probability as the estimate of $S$. And we calculate that
@@
Applying Theorem 3 to Example 2, since $\left|\mathcal{U}\right|\leq|\mathcal{X}|+1$, without loss of generality, we take $\mathcal{U}=\{0,1,2\}$, $u\in\mathcal{U}$, and let the joint distribution of $U$ and $X$ be $p_{jk}=p(u=j,x=k),j=0,1,2,k=0,1$ and define $p_{uj}=p(u=j),j=0,1,2$. Next, we have
$I\left(X;Y|U,S\right)=qH(Y|U,S=1)=q\Big{(}p_{u0}{\rm{H}}_{2}\big{(}\frac{p_{00}}{p_{u0}}\big{)}+p_{u1}{\rm{H}}_{2}\big{(}\frac{p_{10}}{p_{u1}}+p_{u2}{\rm{H}}_{2}\big{(}\frac{p_{20}}{p_{u2}}\big{)}\Big{)}.$
According to $\hat{s}^{*}(u,z)$ in Theorem 3, for $z=0$, we calculate that
@@
$C(D)$ $\geq\max_{p_{UX}}\{\min\{q{\rm{H}}_{2}(p),I\left(U;Z\right)+I\left(X;Y|U,S\right)\}:\sum_{i=0}^{1}\sum_{j=0}^{2}p(z=i,u=j)d_{ij}\leq D\}$
@@
Applying Corollary 2 to Example 2, we get the optimal estimator for $z=0$ and $z=1$,
@@
where we used the independence of $S$ and $X$. Thus, the expected distortion of the optimal estimator is $D=\sum_{x}P_{X}(x)c(x)=P_{X}(0)c(0)=p\min\{q,1-q\},$ Therefore, we obtain
$C(D)$ $=\max_{p\in[0,1]}\{\min\{q{\rm{H}}_{2}(p),I(X;Z)\}:p\min\{q,1-q\}\leq D\}.$
F-B Two upper bounds to Example 2